OpenStudy (anonymous):

find a(n) for the arithmetic sequence a(12)=7,a(14)=15

OpenStudy (anonymous):

a(n) = a(n-1) + d (15 = a(14) + d) subtract (7 = a(12) + d) 8 = 2a, divide by 2, a = 4 now you have y = 2(n-1) + d, and 7 = 2(12) + d, --> 7 = 24 + d, subtract 24, d = -17 So now you have a = 2, d = -17 a(n) = 2(n-1) - 17